### v1

# Evaluating Grounded Reasoning by Code-Assisted LLMs for Mathematics

Abstract
Assisting LLMs with code generation improved their performance on mathematical reasoning tasks. However, the evaluation of code-assisted LLMs is generally restricted to execution correctness, lacking a rigorous evaluation of their generated programs. In this work, we bridge this gap by conducting an in-depth analysis of code-assisted LLMs generated programs in response to math reasoning tasks. Our evaluation focuses on the extent to which LLMs ground their programs to math rules, and how that affects their end performance. For this purpose, we assess the generations of five different LLMs, on two different math datasets, both manually and automatically. Our results reveal that the distribution of grounding depends on LLMs’ capabilities and the difficulty of math problems. Furthermore, mathematical grounding is more effective for close-source models, while open-source models fail to employ math rules in their solutions correctly. On MATH500, the percentage of grounded programs decreased to half, while the ungrounded generations doubled in comparison to ASDiv grade-school problems. Our work highlights the need for in-depth evaluation beyond execution accuracy metrics, toward a better understanding of code-assisted LLMs capabilities and limits in the math domain.
Evaluating Grounded Reasoning by Code-Assisted LLMs for Mathematics
Zena Al-Khalili    Nick Howell    Dietrich Klakow Saarland Informatics Campus, Saarland University, Germany {zakhalili,nhowell,dietrich.klakow}@lsv.uni-saarland.de
1 Introduction
Large Language Models (LLMs) have recently achieved outstanding performance on complex reasoning tasks such as mathematical reasoning, powered by scale and multi-step reasoning approaches, particularly Chain-of-Though (CoT)Wei et al. (2022), that require an LLM to generate the explicit reasoning steps, before generating the final answer.
Despite its success, investigating CoT reasoning steps revealed critical flows of LLMs, such as committing calculation errors Gao et al. (2023) and generating false positive chains Lyu et al. (2023), i.e: containing reasoning errors yet generating correct final answers.
Code-assisted reasoning approaches Gao et al. (2023); Chen et al. (2022); Lyu et al. (2023); Gou et al. (2023); Mickus et al. (2024); Das et al. (2024) proposed to solve these problems by instructing LLMs to generate programmatic reasoning steps, e.g, Python programs, and delegate the execution of these programs to an interpreter, which avoids calculation errors and ensures faithfulness. Such approaches have been found to further improve LLMs performance on math reasoning tasks.
However, performance improvement is predominantly evaluated by the correctness of the execution outcome Gao et al. (2023); Chen et al. (2022); Gou et al. (2023), rather than the quality of the generated programs and how it affects the reasoning process.
This is problematic, as the generated programs can be relaying on exhaustive searches or memorized information to generate correct answers instead of grounding the reasoning process to domain concepts, i.e: mathematical rules, producing less generalizable programs that are harder to verify and trust. Figure 1 show such programs generated by GPT-4 and StarCoder2 when solving math problems from the ASDiv data.
Figure 1: Programs generated by GPT-4 (left) that uses a brute force search to find the answer to the input question, and StarCoder2 (right) that depends on memorized information rather than generating the steps to solve the input question.
The goal addressed in this work is to evaluate the actual reasoning capabilities of code-assisted LLMs when solving mathematical tasks by analyzing their generated intermediate programs.
In this evaluation, we focus mainly on assessing the extent to which code-assisted LLMs ground their reasoning process to domain concepts and math rules when generating their solutions, and how grounding, or its absence, affects end performance on mathematical tasks.
Code enables access to domain concepts such as math rules, via libraries (e.g: Sympy), which can aid LLMs generate programs that are grounded to world knowledge (i.e: mathematical rules), aligning more to human-like reasoning. Given this we expect grounded programs to be trustworthy, easier to verify, and probably with a higher solve rate, while ungrounded programs to be more error-prone, harder to trust, and degrade end-performance.
We investigate the extent to which this is true by seeking answers to the following research questions.
  • RQ1: To what extent do code-assisted LLMs ground their reasoning to relevant mathematical concepts?
  • RQ2: How do code-assisted LLMs implement these concepts?
  • RQ3: What is the relationship between grounding to math concepts and end performance on mathematical tasks?
  • RQ4: How much do the answers to these questions vary with more challenging math problems?
To answer these questions, we start by manually analyzing a subset of intermediate programs, generated by GPT4o-mini, GPT4, Qwen2.5, Llama3, and StarCoder2, for grounding and its implementation style, when solving math problems from the ASDiv and MATH500 datasets (3.3). Given the observations from the manual analysis, we design a taxonomy reflecting the different grounding levels used by evaluated models (3.4). To extend the analysis to the full set of generated programs, we employ two automated evaluation methods: an LLM-Judge using the latest o3-mini model, and our proposed method, Code-Structure Judge, that trains a Decision Tree classifier with features extracted from programs’ Abstract Syntax Tree (3.5). We find Code-Structure judge to achieve higher agreement with human judgment than LLM-judge, with $81\%$ and $73\%$, respectively (5.1); therefore, we employ it for the large-scale evaluation of all generated programs.
To the best of our knowledge this is the first work to analyze generated programs of code-assisted LLMs on math reasoning tasks. We summarize our takeaways from Section (5.2) as following:
  1. 1.
    Code-assisted LLMs ground their programs in math knowledge either by invoking concepts from math library or implementing the concepts from scratch.
  2. 2.
    Code-assisted LLMs also relay on brute-force searches or memorized information to approach some math problems.
  3. 3.
    Grounded generations to math knowledge are more effective for GPT models and Qwen, while open-source models fail to correctly ground their reasoning to useful concepts.
  4. 4.
    The grounding distribution completely changed on the challenging MATH500 data, pushing GPT models and Qwen to double their reliance on brute-force search and memorized information, at the expense of more mathematically-grounded solutions.
Our in-depth evaluation highlights the need for more holistic assessment of LLMs generations, beyond execution accuracy metrics, which don’t provide insightful conclusions about models’ actual capabilities and limits in the math domain.
2 Related Work
Programs as Intermediate Steps for Math Reasoning Code-assisted reasoning approaches such as Chen et al. (2022); Gao et al. (2023); Imani et al. (2023); Lyu et al. (2023); Das et al. (2024) prompt LLMs to generate programs instead of intermediate steps in natural language, which have been found to improve the performance of LLMs on math reasoning tasks. Beyond in-context learning approaches, other work Gou et al. (2023); Mickus et al. (2024) fine-tuned medium-scale language models such as Code-Llama Roziere et al. (2023) on code reasoning paths and achieved comparable performance to closed-source models on math reasoning tasks. However, both in-context and fine-tuned approaches are predominantly evaluated by the correctness of the final answer, overlooking the intermediate programs and how their quality might affect the reasoning process. This work focuses on exploring these programs to accurately evaluate the problem-solving abilities of code-assisted LLMs.
Evaluation of Intermediate Steps This work also relates to several studies that evaluate the intermediate reasoning steps of LLMs in natural language (Golovneva et al., 2022; Hao et al., 2024; Jie et al., 2024; Li et al., 2024) targeting a multitude of evaluation dimensions, such as robustness and faithfulness, or error identification and correction in several reasoning tasks. These works either use a human-written reference chain to compare against the evaluated chain or employ a capable LLM to localize errors and judge the quality of the reasoning chains. Code intermediate steps are unexplored, therefore we aim to analyze the quality of these and how they affect LLMs end performance.
Evaluation of code generated by LLMs. To assess code generated by LLMs, previous work focuses on test-based evaluation and functional correctness, such as (Chen et al., 2021; Liu et al., 2024), others (Ren et al., 2020; Eghbali and Pradel, 2022; Zhou et al., 2023b) proposed metrics to measure how similar a generation is to a reference human-written code, which is expensive to get and doesn’t usually account for generation diversity. (Tong and Zhang, 2024) employed a capable LLM as a judge to localize errors in generated programs on code generation tasks. We draw inspiration from their method to evaluate generated programs, on math reasoning tasks, using an LLM-Judge. Finally, (Dou et al., 2024) analyzed characteristics of LLMs generated programs, on code generation tasks, in terms of code complexity, number of API calls, and types of errors, i.e: bugs, which cause programs to fail. Although insightful, the analysis concluded with a general type of error, commonly shared across different LLMs, namely logic error. In this work, we delve deeper to understand the logic error LLMs commit when writing code to solve a math problem.
3 Evaluation of Intermediate Programs
The main focus of this evaluation is to assess the generated programs of code-assisted LLMs beyond end-to-end evaluation and single-score metrics. We take a step further towards understanding the logic LLMs employ to solve math problems. Therefore, we evaluate their use of math concepts in generated programs. A math concept refers to the principle used to solve a specific mathematical problem, e.g: using the Euclidean Algorithm to find the Greatest Common Divisor of two numbers. Evaluating generated programs for their use of math concepts is analogous to being right for the right reason, as using the math concept in the solution ensures reasoning that is more aligned to human reasoning, and one that can be easily verified and trusted.
3.1 Evaluation Set-up
Given a math problem in natural language, we prompt an LLM to generate a Python program that solves the problem. We discard any generated programs that fail to parse 111 We resolve import errors and global misindentation (where the entire body is misindented). See Appendix A.2 for more details. from our analysis. We provide a list of the most common errors of the generated programs in Appendix D.1
3.2 Evaluation Dataset
We use math problems from two datasets, namely ASDiv Miao et al. (2020) and MATH500 Hendrycks et al. (2021); Lightman et al. (2023). These two sets include diverse problems ranging from grade-school to high-school competition-level questions. We exclude simple problems from the ASDiv data and focus only on more complex skills such as solving a system of equations, finding the greatest common divisor, or the least common multiple. The MATH500 set, on the other hand, has been reported to be more challenging for many LLMs Qwen et al. (2025); Hendrycks et al. (2021), Therefore, we use it to study how LLMs modify their generated programs to tackle more challenging math problems.
3.3 Manual Analysis
Motivated by the lack of fine-grained assessment of LLMs generated code for math reasoning tasks, we conduct a detailed human-manual analysis of the generated programs to answer our first two research questions on whether LLMs ground their generated programs to math concepts and how this grounding is implemented in the generated output. This analysis uses only a subset of the evaluation dataset. sized $300$ programs randomly sampled from the entire set, and consists of two main steps:
Sectioning the programs
We section each program into three blocks: (1) Transcription: This part of the program transcribes the information from the question into variables. (2) Processing: In this section, variables are manipulated via operations and function calls to arrive at the answer. (3) Results collection: Return the final answer of the processing section. Some sections can be fused into one line of code that represents for example processing and returning results at the same step.
Analyzing Processing Sections
We inductively analyze processing code-lines with consideration of the following aspects: logic, implementation style, coherence, verification effort, and trustworthiness.
We check if the code-lines are organized into a coherent sequence of steps that represent a valid implementation of some math concept, that a human can verify against existing possible implementations of the concept.
Additionally, we check whether all steps of a solution are explicitly generated rather than inferred by the model. Inferred steps might involve calculations models can’t compute correctly, due to their proven shortcomings in arithmetic calculations Cobbe et al. (2021); Lewkowycz et al. (2022); Gao et al. (2023), which makes verifying and trusting these solutions harder.
Most importantly, we check how math concepts are implemented in the generated programs. The grounded implementation can happen in several styles, either in terms of primitive operations only, for very simple math problems, and from-scratch implementations, or by calling functions from related math libraries that represent a more abstracted form of the solution. We also observe some implementations relying on brute-force loops that search the space of all possible answers one by one, or lack the processing section altogether.
We organize all these observed trends into six categories that make up our proposed taxonomy of grounding to math concepts and its style.
3.4 Proposed Taxonomy
Given the observations from the manual analysis, we propose to categorize LLMs’ generated programs into six mutually exclusive classes, three of which represent programs grounded to math concepts, in different implementation styles, while the other three represent ungrounded programs that mainly rely on memorized information or exhaustive searches to find the final answer.
  1. 1.
    Conceptual through library calls. Reference a math concept through calls to relevant math libraries, standard, or external.
  2. 2.
    Primitive programs: solutions are expressed in terms of the primitive operations only due to problem simplicity, where no library functionality can be called or implemented.
  3. 3.
    From-scratch Implementation of a library functionality. Instead of a call to a library function, the model implements the same functionality from scratch. The model either inlines this implementation in the generated code or write it as a custom function to be called when required.
  4. 4.
    Brute-Force: The program searches through all possible values to find the answer without guiding the search with any math knowledge.
  5. 5.
    Disorganized: The program consists of incoherent steps that seem to be a mix of the previous classes. Usually include variables defined but not used, or the opposite.
  6. 6.
    No Logic: These programs skip the processing section altogether, merely returning a result without explicitly generating the steps to arrive at it. (Generating the logic as comments in natural language is also considered No Logic.)
3.5 Automated Evaluation
Figure 2: Confusion matrices, showing counts, of Code-Structure judge (left) and LLM-judge (right) on the held-out set of the annotated data from ASDiv. Overall accuracies are $81\%$ and $73\%$ for Code-Structure judge (ours) and LLM-judge (OpenAI o3-mini), respectively. ’FS Imp’: From-scratch implementation, ’BF’: Brute Force.
Extending the human manual analysis to the entire evaluation dataset consumes a lot of time and effort. Therefore, we look into automating the evaluation by training classifiers that label generated programs using one class from our proposed taxonomy. For this purpose, we first annotate a training set using categories from the taxonomy, then we train a decision tree classifier using features extracted from the programs. We compare the trained classifier to an LLM-Judge and evaluate both judges on a held out annotated set. We pick the more accurate judge for the large-scale evaluation.
3.5.1 Training Set Annotation
To train the automated evaluation mehtods, two authors annotate a randomly sampled subset of the generated programs using labels from the taxonomy. The annotated set is $300$ programs from ASDiv split. $210$ examples were used for training, while the rest are held out for measuring the performance of the automated judges.
The annotation guidelines were developed based on observations from the manual analysis. After experimentation with different annotations schemes, we found that human assignment of per-program labels produced low inter-annotator agreement (IAA), while per-line labeling produced very high agreement. We thus selected a per-line annotation scheme, along with a simple algorithm for assigning a program level label based on the per-line labels. More on the annotation guidelines and annotators background can be found in Appendix B.
Double independent annotation was performed on 50 of the 300 samples, with $93\%$ line-granularity and $100\%$ sample-granularity agreement. Disagreements were then adjudicated, and guidelines were updated to resolve the ambiguities.
3.5.2 Automated Evaluation Methods
Code-Structure Judge: Decision Tree Model
We note that each class of the taxonomy tends to rely on specific characteristics of the code structure. Hence, we propose using features from the code structure to train a decision tree classifier that will be used to evaluate generated programs. The features to train the classifier are extracted from the Abstract Syntax Tree (AST) of each program and are the following:
  • Number of function calls, including calls to functions that model writes itself.
  • Number of import statements.
  • Number of built-in operations. e.g: +, < , …
  • Number of control flow statements: e.g, If, for, break, ..
  • Number of variables defined but not used. Number of variables used but not defined.
The max depth used for the decision tree model is $5$. We experimented with other models, such as SVM and Random Forest, but found no further gains. A neural classifier, on the other hand, would require much more training data and, consequently, more annotation and human effort.
LLM-Judge
Following other related work that utilize LLMs as judge Tong and Zhang (2024), we employ an LLM to assess a given program and assign a taxonomy class to that program. For this purpose, we provide a detailed description of the taxonomy classes in the prompt, and ask the LLM-judge to carefully analyze the input program, finally, we ask it to provide a single class number that fit the program most. The prompt includes no reference programs of taxonomy classes. The full prompt is provided in Appendix A.4.
We experimented with two models from OpenAI most advanced models, namely GPT4o and O3-mini, and found that O3-mini achieved slightly better results as a judge. We specify the reasoning effort of the model to be high and allow for $20,000$ max output tokens, including reasoning tokens.
4 Experimental Setup
4.1 Evaluated Models
We analyze code generations of several LLMs, including open-source models such as: StarCoder2 15B Lozhkov et al. (2024), Llama 3.1 8B Dubey et al. (2024) and Qwen2.5 7B Qwen et al. (2025). We use the instruction-tuned version of these models. Furthermore, we evaluate GPT-4 Achiam et al. (2023) and GPT4o-mini from OpenAI. All models were evaluated with greedy decoding. Implementation details are in Appendix A.1
4.2 Prompts
We adept the prompt from Gao et al. (2023) to evaluate LLMs on the ASDiv dataset, with only three demonstrations rather than eight, as no further gain was observed with the full set of demonstrations. For the MATH500 data, we prompt LLMs using demonstrations from the training split of the dataset, adapted from Gou et al. (2023). We evaluate GPT4o-mini and Qwen in zero-shot settings, as we observe models tend to generate less generations in natural language using this setting. The two full prompts can be found in Appendix A.3.
Model Conceptual Primitive FS Implementation Brute-Force Disorganized No Logic
GPT4o-mini $90.5\%$ $96.5\%$ $86.3\%$ $77.7\%$ $93.7\%$ $85.7\%$
GPT4 $86\%$ $86\%$ $100\%$ $68.4\%$ $60\%$ $63.6\%$
Qwen2.5 $91.5\%$ $78.3\%$ $89.4\%$ $100\%$ $60\%$ $100\%$
StarCoder2 $48\%$ $44.4\%$ $76.4\%$ $47.6\%$ $30.7\%$ $61.5\%$
Llama3.1 $35.7\%$ $48.1\%$ $37.5\%$ $53.7\%$ $33.3\%$ $57.5\%$
Table 1: Execution (macro) accuracy in percentages per taxonomy class for all evaluated models on the ASDiv math problems. ’FS’ is From Scratch.
5 Results and Analysis
5.1 Comparison of different automated methods
To compare automated evaluation methods, we measure the agreement to human judgment using accuracy to ground truth labels assigned by annotators to each generated program.
Code-Structure Judge achieved a mean accuracy of $81\%$ with a standard deviation of $0.005$ over four different seeds. Figure 2 (left) shows the confusion matrix of the Code-Structure judge on the held out set from ASDiv data.
We notice that Code-Structure judge achieves both high precision and recall on most of the classes, except for the Disorganized class with a low recall of $0.16$. The Disorganized programs can be a mix of other classes and the distinction between these can be semantic rather than structural. Class Primitives also has a lower precision than other classes with $0.72$. Appendix C.2 include some incorrectly classified programs by the Code-Structure judge, and the full table of precision, recall and F1-score.
LLM-judge, on the other hand, achieved lower accuracy with only $73\%$. The Primitives class has a low precision of $0.56$, while the Disorganized and No Logic classes have much lower recall of $0.25$ and $0.37$ respectively, decreasing the overall accuracy of this judge by $8\%$ in comparison to the Code-Structure judge. Figure 2 (right) shows the confusion matrix of LLM-judge in comparison to ground truth labels.
We conducted a qualitative error analysis of some incorrectly classified instances of the No Logic class and found that the LLM-judge tends to mistake the step of transcribing information from the question to be part of the logic, and consequently classifies the instance as Primitive instead.
5.2 Insights on code-assisted LLMs from evaluating their intermediate programs
We employ the Code-Structure judge for large-scale evaluation of the entire set of generated programs in response to math word problems, aiming to answer our research questions:
Do code-assisted LLMs ground their generations to math rules?
The distribution of taxonomy classes per model on ASDiv is presented in Figure 3. We observe that while all models generate mathematically-grounded programs, there are differences in the way they ground: GPT models and Qwen mainly ground through library calls, while the other open-source models ground through primitives. All models depend less on from-scratch implementation. Older models (StarCoder2, Llama3.1, and GPT4) generate more ungrounded programs (29%, 43%, and 26%, respectively) as compared with the newer models (GPT4o-mini 14% and Qwen 7%). The majority GPT4 and Llama3.1 ungrounded generations are “Brute-force search” (20% and 26% of all generated programs, respectively), while StarCoder2 produces more “No logic” solutions (16% of all generations). All models generate very small percentage of disorganized programs.
Figure 3: Distribution of taxonomy classes for all evaluated models in percentages on ASDiv problems. Classes are: ’Con’: Conceptual, ’Prim’: Primitives, ’FS Imp’: From scratch Implementation, ’BF’: Brute-Force, ’D’: Disorganized, and ’NoL’: No Logic.
What is the relation between grounding generations to math rules and end performance?
Table 1 provide the execution macro accuracy per taxonomy class for all evaluated models on the ASDiv math problems. We note that GPT models and Qwen grounded programs achieve higher accuracy than the grounded programs generated by less capable open-source models, indicating that grounding is not as effective for these models. On the other hand, many models perform well on the ungrounded programs; both Llama3.1 and StarCoder2 achieve higher accuracy on No Logic than Conceptual programs. This implies that these models attempt to solve math problems by memorizing the answers without solving the underlying reasoning problems. Similarly, many of the models employ Brute-Force searches to find the final answer, without guiding the search with any mathematical rules. Studying the correlation between the execution accuracy and grounding provided no useful insights, as models seems to perform well on all categories. [^2] Given the small sample size of ungrounded generations of GPT models and Qwen, we can’t conclude much about their performance on these categories. Overall, we can see that grounding generations to math knowledge is more effective for GPT models and Qwen. Conversely, open-source models depend on other hacks to approach math problems, which might indicate their inability to perform a successful grounding.
Model Conceptual Primitive FS Implementation Brute-Force Disorganized No Logic
GPT4o-mini $56.5\%$ $59.2\%$ $48.7\%$ $48.6\%$ $64.2\%$ $56.8\%$
Qwen2.5 $48.6\%$ $54.2\%$ $54.5\%$ $60.2\%$ $23.5\%$ $62.5\%$
StarCoder2 $10\%$ $17.1\%$ $0.0\%$ $25.0\%$ $8.3\%$ $46.1\%$
Table 2: Execution (macro) accuracy in percentages per taxonomy class for three evaluated models on the MATH500 problems. ’FS’ is From Scratch.
Why do LLMs perform differently within a class?
GPT4o-mini has very high accuracy on the Disorganized class, unlike all other models. To further understand this counter-intuitive phenomenon, we analyze the $16$ generations in this class by GPT4o-mini and find that out of all generations, $14$ are false positives, with $9$ Primitives and $5$ No Logic. We also investigate the reason why Llama3.1 and StarCoder2 generally have degraded performance on the grounded classes. On the Conceptual class, we qualitatively analyze 20 programs by the two models and find that 18/20 programs invoke the wrong concept, indicating a poor ability to invoke relevant math concepts. The other 2/20 programs incorrectly transcribe information from the question. Furthermore, on the Primitive class, we conduct error analysis on 20 programs generated by the two models and find that 9/20 were false positives of Disorganized and No Logic, and 8/20 have inferred step error, where the model infer the calculation of one step rather than generating the steps to calculate it, causing the error to propagate through the rest of the code. Additionally, 2/20 incorrectly transcribed information from the question and 1/20 had wrong ground truth answer. Finally, on the class From-scratch Implementation, Llama3.1 has the lowest accuracy in comparison to other models. Examining 10 random generations from this class, we find that the model implements a different solution than what the problem requires. For example implementing the code for the Greatest Common Divisor problem, when the question is asking for the Least Common Multiple or vice versa. Starcoder2, on the other hand, successfully implements many solutions.
How do more challenging math problems affect LLMs’ grounding?
To answer this question, we evaluate programs generated by GPT4o-mini, Qwen2.5 and StarCoder2 in response to problems from the MATH500 dataset, which has been found to be more challenging for LLMs Hendrycks et al. (2021). Figure 4 illustrates the distribution of taxonomy classes per evaluated LLM on MATH500.
Figure 4: Distribution of taxonomy classes for all evaluated models in percentages on MATH500 dataset. Classes are: ’Con’: Conceptual, ’Prim’: Primitives, ’FS Imp’: From scratch Implementation, ’BF’: Brute-Force, ’D’: Disorganized and ’NoL’: No Logic.
We notice that more challenging data changes the distribution of grounded generations. The number of conceptual programs decreased to half for all models, while the number of ungrounded programs, specifically Brute-Force and No Logic, doubled for all models. This indicates that the difficulty of the problem might push models to rely less on grounded program.
As the number of programs in the No Logic class increased, we qualitatively analyze some generations in this class and find that for StarCoder2 and Qwen2.5 many instances actually just return the answer right away, while for GPT-4o-mini, many instances solve the questions with comments in natural language, rather than programmatic steps that can be executed to get the answer. We still consider these instances to be No Logic as these instances resemble reasoning in natural language, and code is not efficiently helping in any way in finding the correct answer. The preference of textual reasoning over code might be due to problem complexity, as empirically investigated and discussed in (Chen et al., 2025), where they show that some GPT models prefer textual reasoning over code reasoning depending on the complexity of the task.
As for the relation between execution accuracy and grounding, Table 2 presents execution macro accuracy per class per model on the MATH500 problems. We observe that the accuracy on ungrounded programs is higher or on par with accuracy on grounded generations, which is more pronounced for both open-source models  [^3] The high accuracy on Disorganized class for GPT4o-mini and Qwen is due to false positives.. This is problematic, as relying on memorized information or brute-force searches to find the answer is not trustworthy, and can’t provide consistent performance. We also study the library calls made by the evaluated LLMs on the two datasets and find that GPT4o-mini and Qwen exploit a greater variety of library calls in both datasets, but the distribution of those library calls shifts from ASDiv (where math.gcd, abs, math.lcm, min are most frequent) to MATH500 (where sympy.solve dominates). Starcoder2, on the other hand, doesn’t employ many libraries and barely generates Conceptual programs on the MATH500 dataset.
Finally, we look at the relation between math subjects and taxonomy classes to check if models generate a certain type of programs depending on the subject of the given problem. Both GPT4o-mini and Qwen2.5 tend to use more Primitive solutions for easier problems, such as Prealgebra and Precalculus. Additionally, they employ Conceptual programs for Algebra and Probability problems. Notably, both models also exploit memorized information or brute-force search for Intermediate Algebra, Number Theory, and some of the Algebra problems. Figure 9 in Appendix D.2 illustrates the distribution of taxonomy classes per math subject from the MATH500 dataset. StarCoder2, on the other hand, heavily depends on the Primitive solutions for most of the subjects, demonstrating a lack of diversity in the logic it employs for different types of math problems.
6 Conclusion
In this work, we conducted an in-depth analysis of code-assisted LLMs generated programs in response to math problems. Our assessment focuses on the extent to which LLMs employ math concepts in the generated programs and how this affects their end performance. We categorized the generated programs into six different categories, which make our proposed taxonomy of grounding to math concepts. We annotated a corpus of programs using classes from the taxonomy and trained a decision-tree model that we call Code-Structure Judge. The Code-Structure judge outperformed an LLM-judge baseline and was employed for a large-scale evaluation of more generated programs. Our findings show that models employ grounded reasoning, as well as ungrounded reasoning, comprising the use of brute-force search and memorized information. The use of ungrounded hacks depends highly on LLMs capabilities and problems difficulty. Our work emphasizes the need for holistic evaluation of code-assisted LLMs, and we hope our findings will inspire future work to further study why LLMs employ ungrounded solutions and how to mitigate this phenomenon.
Limitations
We note there’s limitations to our work. First: Code-structure Judge is still not accurate on the Disorganized class, causing many false positives. Second, we depend on one prompting technique and don’t compare performance when utilizing prompts to improve generated programs with further refinement, such as Self-Refine (Madaan et al., 2024) or Code-based Self-Verification (Zhou et al., 2023a).
Acknowledgments
The authors would like to thank Vagrant Gautam for their mentorship and useful feedback. We are also grateful to Marius Mosbach and Ellie Pavlick for feedback during the early stages of this work. We also thank Miaoran Zhang for the useful discussions and Tural Mammadov for help in the experimental infrastructure. The authors received funding from the DFG (German Research Foundation) under project 232722074, SFB 1102.
Appendix A Experimental Details
A.1 Implementation Details
We use starcoder2-15b-instruct, Meta-Llama-3-8B-Instruct, Qwen2.5-7B-Instruct from HuggingFace Wolf et al. (2020). For OpenAI models we use gpt-4-32K and gpt-4o-mini We use int8bit model quantization for all models except OpenAI models as we observe no major differences in the execution accuracy per model. Finally, we use NVIDIA A100 GPUs 40GB, with batch size = 32 for evaluating Llama3 and StarCoder2 and Qwen2.5 on the ASDiv, which took around one hour. While evaluating these models on the MATH subset took about 3 hours with batch size=16.
A.2 Resolving some execution errors
To resolve import errors we used autoimport (lyz@riseup.net, ). For indentation errors we used the inspect module from the Python standard library as a post processing step.
A.3 PAL Prompt and MATH Prompt
In our experiments, we employ PAL prompt from (Gao et al., 2023) to evaluate LLMs on the ASDiv data, however we only include three demonstrations out of eight, as we observe no further gains from including the entire set. The full prompt is in Figure 5.
Figure 5: PAL prompt used to prompt evaluated models with three demonstrations on the ASDiv dataset.
System: Write python program to solve math problems.
User: # Q: Michael had 58 golf balls. On tuesday, he lost 23 golf balls. On wednesday, he lost 2 more. How many golf balls did he have at the end of wednesday?
Assistant: # solution in Python: def solution(): """Michaelhad58golfballs.Ontuesday,helost23golfballs. ␣␣␣␣␣␣␣␣Onwednesday,helost2more.Howmanygolfballsdidhehaveat ␣␣␣␣␣␣␣␣theendofwednesday?""" golf_balls_initial = 58 golf_balls_lost_tuesday = 23 golf_balls_lost_wednesday = 2 golf_balls_left = golf_balls_initial \ - golf_balls_lost_tuesday \ - golf_balls_lost_wednesday result = golf_balls_left return result
User: # Q: Olivia has \$23. She bought five bagels for \$3 each. How much money does she have left?
Assistant: # solution in Python: def solution(): """Oliviahas\$23.Sheboughtfivebagelsfor\$3each.Howmuch ␣␣␣␣␣␣␣␣moneydoesshehaveleft?""" money_initial = 23 bagels = 5 bagel_cost = 3 money_spent = bagels * bagel_cost money_left = money_initial - money_spent result = money_left return result
User: # Q: There were nine computers in the server room. Five more computers were installed each day, from monday to thursday. How many computers are now in the server room?
Assistant: # solution in Python: def solution(): """Therewereninecomputersintheserverroom.Fivemore ␣␣␣␣␣␣␣␣computerswereinstalledeachday,frommondaytothursday.How ␣␣␣␣␣␣␣␣manycomputersarenowintheserverroom?""" computers_initial = 9 computers_per_day = 5 num_days = 4 # 4 days between monday and thursday computers_added = computers_per_day * num_days computers_total = computers_initial + computers_added result = computers_total return result
Th prompt employed for evaluating LLMs on the MATH data is found in Figure6
Figure 6: Prompt structure for evaluating on MATH dataset.
System: Write python programs to solve math problems
User: # Q: Find the coefficient of $x^{3}$ when $3(x^{2}-x^{3}+x)+3(x+2x^{3}-3x^{2}+3x^{5}+x^{3})-5(1+x-4x^{3}-x^{2})$ is simplifie.
Assistant: # solution in Python: from sympy import symbols, simplify def solution(): x = symbols(’x’) expr = 3*(x**2 - x**3 + x) + 3*(x + 2*x**3 - 3*x**2 + 3*x**5 + x**3) - 5*(1 + x - 4*x**3 - x**2) simplified_expr = simplify(expr) x3_coefficient = simplified_expr.as_coefficients_dict()[x**3] result = x3_coefficient return result
User: # Q: The surface area of a sphere with radius r is $4\pi r^{2}$. Including the area of its circular base, what is the total surface area of a hemisphere with radius 6 cm? Express your answer in terms of pi
Assistant: # solution in Python: import math def solution(): radius = 6 # Surface area of the hemisphere hemisphere_area = 2 * math.pi * radius**2 # Area of the circular base base_area = math.pi * radius**2 # Total surface area total_surface_area = hemisphere_area + base_area result = total_surface_area / math.pi return result
User: # Q: Monica tosses a fair 6-sided die. If the roll is a prime number, then she wins that amount of dollars (so that, for example, if she rolls 3, then she wins 3 dollars). If the roll is composite, she wins nothing. Otherwise, she loses 3 dollars. What is the expected value of her winnings on one die toss? Express your answer as a dollar value to the nearest cent.
Assistant: # solution in Python: def solution(): # Probabilities of each outcome prime_prob = 1 / 6 composite_prob = 1 / 3 otherwise_prob = 1 / 6 # Expected value of each outcome prime_expected_value = (2 * prime_prob) + (3 * prime_prob) + (5 * prime_prob) composite_expected_value = 0 * composite_prob otherwise_expected_value = -3 * otherwise_prob # Total expected value total_expected_value = prime_expected_value + composite_expected_value + otherwise_expected_value # Dollar value to the nearest cent result = "{:.2f}".format(total_expected_value) return result
A.4 LLM-Judge Prompt
Figure 7 present the prompt used to prompt o3-mini as an LLM-Judge
Figure 7: The prompt for the LLM-Judge task with o3-mini
User:
You are an expert code judge that analyze code based on the
following classes:
 1- Conceptual through library calls. Reference a math concept
 through calls to relevant math libraries, standard or external.
 2- primitive solution: programs are expressed in terms of the
 primitive operations only due to problem simplicity, where no
 library functionality can be called or implemented.
 3- From-scratch Implementation of a library functionality.
 Implements a library functionality from scratch. implementation
 is inlined in the generated code, or can be a custom function
 to be called when required.
 4- Brute-Force. The program search through all possible values
 to find the answer without guiding the search with some math
 knowledge.
 5- Disorganized: the program consists of incoherent steps that
 seem to be a mix of the previous classes. Usually include
 variables used but not defined or the opposite.
 6- No Logic: These programs merely return a result without
 explicitly generating the steps to arrive at it, transcribing
 information from the question only without further processing
 the information is also No logic. generating the logic as
 comments doesn’t count either.
Instructions:
- given an input program your task is to analyze it and then
provide a class number from the list above.
- don’t fix the code.
[…]
Appendix B Training Set Annotation Guidelines and Annotators Background
B.1 Annotation format
Each line of program code is prefixed with five characters of the form TPIR plus space.
Columns:
  • T.
    transcription
  • P1-5.
    processing (blank implies "no logic")
  • I.
    inference-time computation
  • R.
    result collection
B.2 Transcription
Purely transcriptive statements that transcribe either data (e.g. numbers) or relations between data (e.g. equations) from the question.
Purely transcriptive statements can still contain operations that are explicitly described in the problem statement; these operations do not count as processing.
     """ One number is twice as
     large as another. The smaller
     number is 3. Find the other
     number """
     def solution():
T      x = 3
T  R   y = 2*x
   R   return y
Even if the model uses the question to compute at inference time, without pure transcription of data or relations no T is marked:
     One number is twice as large
     as another. The sum of the
     numbers is 12. Find the gcd of
     the two numbers
     def solution():
  I    x = 4
  I    y = 8
 1 R   z = math.gcd(x, y)
   R   return z
If the model has performed partial computations (collapsing e.g. a pure transcription and true processing) then it receives no T; it can possibly receive other labels.
     Two boards have total length
     10; the long board is 2 longer
     than the short board. Find the
     lengths of the two boards
     def solution():
T      total_length = 10
T      difference = 2
 2IR   short_length = (total_length -
         difference) / 2
T  R   long_length = short_length +
         difference
   R   return short_length, long_length
B.3 Processing types
Comments are never marked as anything!!
  • 1.
    conceptual lib
    • calls a library function later
    • mark the entire function; but not the comments
  • 2.
    conceptual primitive
    • uses primitives in a way that is correct, and cannot be simplified by lib
  • 3.
    from scratch implementation
    • plausibly correct, looks like inlined implementation of a lib function
    • writes a function for itself, then calls (mark the call as implementation as well.)
  • 4.
    brute-force
    • search through all space
    • mark the entire loop
  • 5.
    disorganised
    • probably incorrect (more or less random?, disorganised)
Empty ’processing type’ field (i.e., a space " ") indicates that no processing occurs.
B.4 Inference-time computation
If the model skips steps (either entirely, or in part (in which case there will still be a processing label)) Then the line gets the "inference-time computation" flag
In the case of the model entirely precomputing, the processing type will probably be empty and the I flag will appear after.
     def solution():
  I    x = 6 # lcm of 3 and 2
       return x
If the model has performed symbolic manipulation to avoid the use of symbolic equation libraries, it receives I as processing/transcription labels.
B.5 Result collection
Lines that return the final answer, or that store the variable holding the final answer, are marked R in the final column.
     def solution():
  IR   x = 6 # lcm of 3 and 2
   R   return x
     """ One number is twice as
     large as another. The sum of
     the numbers is 12.  Find the
     gcd of the two numbers"""
     def solution():
  I    x = 4
  I    y = 8
 1 R   z = math.gcd(x, y)
   R   return z
B.6 Annotators Background
The annotation process of the generated programs is time-consuming and isn’t feasible to do at a larger scale because it requires expertise with code understanding. Both annotators are experts in computer science and have done prior reading on programming languages and related topics, that make them a better fit for the annotation process.
Furthermore, programmatic steps and code structure patterns are less subjective than natural language, carrying less nuance, which can lead to disagreement or bias. For example, it is hard to mistake a library call from a brute-force program.
Appendix C More Results on Code-Structure Judge and LLM-Judge Performance
C.1 Per-class $F_{1}$ scores
Table 3 provides precision, recall, and F1 scores of both code judges.
Precision Recall $\mathbf{F_{1}}$
Conceptual $1$ $1$ $1$
Primitive $0.72$ $0.95$ $0.82$
From-scratch Implementation $0.66$ $1$ $0.8$
Brute-Force $0.88$ $1$ $0.94$
Disorganized $0.5$ $0.16$ $0.25$
No Logic $0.91$ $0.68$ $0.78$
Precision Recall $\mathbf{F_{1}}$
Conceptual $1$ $0.88$ $0.94$
Primitive $0.56$ $0.95$ $0.71$
From-scratch Implementation $0.83$ $0.83$ $0.83$
Brute-Force $0.83$ $0.93$ $0.88$
Disorganized $0.6$ $0.25$ $0.35$
No Logic $0.75$ $0.37$ $0.5$
Table 3: Precision, recall, and $F_{1}$ of the Code-Structure Judge (top) and LLM-Judge (bottom)
C.2 Examples of incorrectly classified programs
We provide a few examples of programs where Code-Structure judge misclassifies Disorganized programs and their true label in Figure C.2.
def find_other_number(): difference = 100 one_number = 91 other_number = one_number + difference return other_number result = find_other_number() True label: “Primitive”
def smallest_page_number(): # The least common multiple (LCM) of 4 and 13 lcm = 4 * 13 # Since 4 and 13 are co-prime return lcm result = smallest_page_number() True label: “No Logic”
Figure 8: Instances that were labeled “Disorganized” by the Code-Structure Judge, and their true label.
Appendix D Further Analysis of Generated Programs
D.1 Most common bugs in the generated programs on MATH500 dataset
StarCoder2 generated 60 programs with bugs: 23 of them with undefined symbols and 15 that were calling undefined functionality from libraries. e.g, calling ’log3’ from the math library. Qwen2.5 generated 40 programs with bugs as follows: 8 were with undefined symbols, 7 with undefined attributes of objects e.g, a float object has no attribute denominator. GPT4o-mini generated only 5 programs with bugs, two of which have undefined symbols.
D.2 Analysis of programs per math subject
Figure 9 illustrates the distribution of taxonomy classes per math subject from the MATH500 dataset.
(a) GPT4o-mini
(b) Qwen2.5
(c) StarCoder2
Figure 9: Distribution of taxonomy classes counts per math subject from the MATH500 dataset for evaluated models. Classes: ’FS Imp’: From-Scratch Implementation, ’BF’: Brute-Force ’Disorg’: Disorganized.